HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

COUPLED METHYL INTERNAL ROTATIONS IN HETEROCYLIC MOLECULES: THE MICROWAVE SPECTRA OF 2,4-DIMETHYLPYRROLE AND 2,3-DIMETHYLFURAN

\underline{M. G. BARTH}^{1}, M. SCHWELL^{1}, I. KLEINER^{2}, G. BRANNYS^{3}, H. V. L.

NGUYEN^{1,4},^{1}Univ Paris Est Creteil and Université Paris Cité, CNRS, LISA, F-94010 Créteil, France,^{2}Université Paris Cité and Univ Paris Est Creteil, CNRS,

LISA, F-75013 Paris, France,^{3} Institute of Physical Chemistry, RWTH Aachen University, Landoltweg 2, D-52074 Aachen, Germany,^{4} Institut Universitaire de

France (IUF), F-75231 Paris, France

The rotational spectra of the three isomers of methylpyrrole^{1}^{,}^{2}^{,}^{3} and 2,5-dimethyl-pyrrole^{4} as well as those of the two isomers of methylfuran^{5}^{,}^{6} and 2,5-dimethylfuran^{7} have been investigated. In both series, the effects of methyl internal rotations were observed. The barrier hindering the methyl torsion are mainly impacted by the positions of the methyl groups on the aromatic ring, which do not only reflect the steric but also electronic properties of the molecules.

In the present work, we studied two further cases with two methyl tops, 2,3dimethylfuran and 2,4-dimethylpyrrole. The spectra were measured using a molecular jet Fourier transform microwave spectrometer operating from 2 to 26.5 GHz. The barriers to internal rotation were determined for both molecules. The hyperfine splittings due to the quadrupole coupling of the^{14}N nucleus in 2,4-dimethylpyrrole were fully resolved. The quadrupole coupling constants χ_{aa}, χ_{bb}, and χ_{cc} were determined with high accuracy and calculated well using Bailey's method.^{8} The obtained barriers were finally compared to those found for other molecules in the series to observe a trend for methyl internal rotational barriers in pyrrole and furan derivatives.

1doi:10.1016/S0022-2860(02)00082-0, J. Makarewicz, S. Huber, B. Brupbacher-Gatehouse, A.

Bauder, J. Mol. Struct. 612, 117 (2002).

2doi:10.1080/00268976.2019.1668572, T. Nguyen, C. Dindic, W. Stahl, H.V.L. Nguyen, I. Kleiner, Mol. Phys. 118, 1668572 (2020).

3doi:10.1016/j.jms.2020.111351, T. Nguyen, W. Stahl, H.V.L. Nguyen, I. Kleiner, J. Mol. Spectrosc.

372, 111351 (2020).

4doi:10.1063/5.0049418, T. Nguyen, W. Stahl, H.V.L. Nguyen, I. Kleiner, J. Chem. Phys. 154, 204304 (2021).

5doi:10.1063/1.1671739, W.G. Norris, L.C. Krisher, J. Chem. Phys. 51, 403 (1969).

6doi:10.1246/bcsj.44.2344, T. Ogata, K. Kozima, Bull. Chem. Soc. Jpn. 44, 2344-2346 (1971).

7doi:10.1016/j.jms.2017.11.007, V. Van, J. Bruckhuisen, W. Stahl, V. Ilyushin, H.V.L. Nguyen, J.

Mol. Spectrosc. 343, 121-125 (2018).

8doi:10.1016/S0301-0104(99)00342-0, W. C. Bailey, Chem. Phys. 252, 57 (2000).

p-number: p154  Submitted on Wed, 14 Jun 2023 20:50:58 +0200